# Individualized and Interpretable Sleep Forecasting via a two-stage Adaptive Spatial-Temporal Model

### v1

@@ V-D Explainable model for sleep forecast
The results reveal distinct patterns of feature importance across participants, demonstrating both individual variability and common influential factors in sleep quality prediction.
@@ V-D1 Clinical Implications of Feature Importance
The SHAP analysis provides several clinically relevant insights:
@@ V-D2 Model Behavior Validation
    Bidirectional Effects: The presence of both positive and negative SHAP values for different features indicates the model captures complex, relationships. For instance, while increased deep sleep positively influences predictions, excessive restlessness negatively impacts them, reflecting real-world sleep dynamics.
@@ VI Conclusions and Limitations
Our model could inform users and clinicians about future potential deterioration, thereby facilitating individualized and timely interventions. The model’s ability to maintain performance across extended periods (up to 170 days) and diverse sleep patterns validates its potential for real-world deployment in continuous sleep monitoring systems. Through SHAP analysis, we identified Deep Sleep Seconds as the most influential predictor across participants, while revealing significant individual variations in feature importance. This interpretability framework transforms our model from a black box into a transparent tool that provides actionable insights for personalized sleep interventions.
@@
ACKNOWLEDGMENT
The work is supported by the Healthy Living as a Service project (with project number 13948 of the research programme KIC which is (partly) financed by the Dutch Research Council (NWO)). We would also like to thank the facilities, like the Habrok cluster provided by CIT of the University of Groningen, and other support.

### v2

@@ V-D Explainable model for sleep forecast
The results reveal distinct patterns of feature importance across participants, demonstrating both individual variability and common influential factors in sleep quality prediction, as shown in Figure 6.
@@ V-D1 Clinical Implications of Feature Importance
The analysis provides several clinically relevant insights:
@@ V-D2 Model Behavior Validation
    Bidirectional Effects: The presence of both positive and negative values for different features indicates the model captures complex, relationships. For instance, while increased deep sleep positively influences predictions, excessive restlessness negatively impacts them, reflecting real-world sleep dynamics.
@@ VI Conclusions and Limitations
Our model could inform users and clinicians about future potential deterioration, thereby facilitating individualized and timely interventions. The model’s ability to maintain performance across extended periods (up to 170 days) and diverse sleep patterns validates its potential for real-world deployment in continuous sleep monitoring systems. Through analysis, we identified Deep Sleep Seconds as the most influential predictor across participants, while revealing significant individual variations in feature importance. This interpretability framework transforms our model from a black box into a transparent tool that provides actionable insights for personalized sleep interventions.